OpenStudy (zyi6):

Can someone help me with these Algebra 1 questions? Will medal!

OpenStudy (zyi6):

Part A: Use the properties of exponents to explain why 64 raised to the power of 1 over 2 is called the square root of 64. (5 points) Part B: The length of a rectangle is 7 units and its width is square root of 7 unit. Is the area of the rectangle rational or irrational? Justify your answer. (5 points)

OpenStudy (zyi6):

@nincompoop ? could you PLEASE help me? I need it! These are questions off of the study guide, not the acutal test.

TheSmartOne (thesmartone):

Part A: \(\sf \LARGE x^{\frac{a}{b}} = \sqrt[b]{x^a}\) Use that property of exponents to explain it. Part B: Do you know what a rational and irrational number is? And to get the area of a rectangle it is \(\sf Area = Length \times Width\)